OpenStudy (anonymous):

is this a circle graph? (-7,-9) radius of 49?

geerky42 (geerky42):

Not enough information

OpenStudy (anonymous):

x squared plus y squared minus 14x -18y plus 81 equals zero

OpenStudy (cwrw238):

yes this is a circle but the radius is 7

OpenStudy (cwrw238):

no hold on - the center is at (7,9)

OpenStudy (cwrw238):

x^2 + y^2 - 14x - 18x + 81 = 0 complete the square: (x - 7)^2 - 49 + (y - 9)^2 - 81 + 81 = 0 (x - 7)^2 + (y - 9)^2 = 49 - this is the equation in standard form

OpenStudy (cwrw238):

standard form is (x - a)^2 + (y - b)^2 = r^2 where (a,b) is the centre of circle and r = radius

OpenStudy (cwrw238):

do you want the equation of the circle with centre (-7,-9) and radius 49?

OpenStudy (cwrw238):

that would be (x + 7)^2 + (y + 9)^2 = 49^2

OpenStudy (cwrw238):

to get it in the same form as your equation you need to expand the brackets and rearrange

OpenStudy (anonymous):

am i changing the signs bc i got -7,-9

OpenStudy (anonymous):

with a radius of 49